# Rendering Wireless Environments Useful for Gradient Estimators: A Zero-Order Stochastic Federated Learning Method

### v1

@@ B-B Proof of Lemma 2: Expected Norm Squared of the Estimated Gradient
To bound the norm squared of the one-point gradient estimate, we start by letting $\theta_{k}^{\prime}=\theta_{k}+\gamma_{k}\Phi_{k}\sum_{j=1}^{N}\big{(}\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k}\big{)}$
$\begin{split}&\mathbb{E}[\|g_{k}\|^{2}|\mathcal{H}_{k}]\\ =&\mathbb{E}\Big{[}\|\Phi_{k}\|^{2}\Big{(}\sum_{i=1}^{N}h_{i,k+1}\tilde{f}_{i}\big{(}\theta_{k}^{\prime},S_{i,k+1}\big{)}+n_{i,k+1}\Big{)}^{2}\Big{|}\mathcal{H}_{k}\Big{]}\\ \overset{(a)}{\leq}&\alpha_{3}^{2}N\sum_{i=1}^{N}\mathbb{E}\Big{[}\Big{(}h_{i,k+1}\tilde{f}_{i}\big{(}\theta_{k}^{\prime},S_{i,k+1}\big{)}+n_{i,k+1}\Big{)}^{2}\Big{|}\mathcal{H}_{k}\Big{]}\\ \overset{(b)}{=}&\alpha_{3}^{2}N\sum_{i=1}^{N}\mathbb{E}\Big{[}h_{i,k+1}^{2}\tilde{f}_{i}^{2}\big{(}\theta_{k}^{\prime},S_{i,k+1}\big{)}+n_{i,k+1}^{2}\Big{|}\mathcal{H}_{k}\Big{]}\\ \overset{(c)}{\leq}&\alpha_{3}^{2}N\sum_{i=1}^{N}\mathbb{E}\Big{[}\frac{h_{i,k+1}^{2}}{\sigma_{h}^{4}}\Big{(}\|f_{i}(0,S_{i,k+1})\|+L_{S_{i,k+1}}\|\theta_{k}^{\prime}\|\Big{)}^{2}\Big{|}\mathcal{H}_{k}\Big{]}\\ &+N^{2}\alpha_{3}^{2}\sigma_{n}^{2}\\ \leq&\alpha_{3}^{2}N\sum_{i=1}^{N}\mathbb{E}\Big{[}\frac{h_{i,k+1}^{2}}{\sigma_{h}^{4}}\Big{(}\|f_{i}(0,S_{i,k+1})\|+L_{S_{i,k+1}}\|\theta_{k}\|\\ &\hskip 14.22636pt+L_{S_{i,k+1}}\gamma_{k}\|\Phi_{k}\|\Big{|}\sum_{j=1}^{N}\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k}\Big{|}\Big{)}^{2}\Big{|}\mathcal{H}_{k}\Big{]}+N^{2}\alpha_{3}^{2}\sigma_{n}^{2}\\ \overset{(d)}{\leq}&3\alpha_{3}^{2}N\sum_{i=1}^{N}\mathbb{E}\Big{[}\frac{h_{i,k+1}^{2}}{\sigma_{h}^{4}}\Big{(}\|f_{i}(0,S_{i,k+1})\|^{2}+L_{S_{i,k+1}}^{2}\|\theta_{k}\|^{2}\\ &\hskip 14.22636pt+\alpha_{3}^{2}L_{S_{i,k+1}}^{2}\gamma_{k}^{2}\big{(}\sum_{j=1}^{N}\frac{h_{j,k}}{\sigma_{h}^{2}}+n_{j,k}\big{)}^{2}\Big{)}\Big{|}\mathcal{H}_{k}\Big{]}+N^{2}\alpha_{3}^{2}\sigma_{n}^{2}\\ \overset{(e)}{\leq}&\frac{3N^{2}\alpha_{3}^{2}}{\sigma_{h}^{2}}\Big{(}\mu_{S}+L_{S}\|\theta_{k}\|^{2}\Big{)}+3N\alpha_{3}^{4}L_{S}\gamma_{k}^{2}\sum_{i=1}^{N}\mathbb{E}\Big{[}\frac{h_{i,k+1}^{2}}{\sigma_{h}^{4}}\times\\ &\big{(}\sum_{j\neq i}\frac{h_{j,k}^{2}}{\sigma_{h}^{4}}+\frac{h_{i,k}^{2}}{\sigma_{h}^{4}}+n_{j,k}^{2}+2\sum_{j<l}\frac{h_{j,k}h_{l,k}}{\sigma_{h}^{4}}\big{)}\Big{|}\mathcal{H}_{k}\Big{]}+N^{2}\alpha_{3}^{2}\sigma_{n}^{2}\\ \end{split}$
=$\begin{split}\overset{(f)}{=}&\frac{3N^{2}\alpha_{3}^{2}}{\sigma_{h}^{2}}\Big{(}\mu_{S}+L_{S}\|\theta_{k}\|^{2}\Big{)}\\ &+3N^{2}\alpha_{3}^{4}L_{S}\gamma_{k}^{2}\big{(}\frac{(N-1)\sigma_{h}^{4}}{\sigma_{h}^{8}}+\frac{\sigma_{h}^{4}+2K_{hh}^{2}}{\sigma_{h}^{8}}+\frac{N\sigma_{n}^{2}}{\sigma_{h}^{2}}+0\big{)}\\ &+N^{2}\alpha_{3}^{2}\sigma_{n}^{2}\\ :=&c_{2},\end{split}$ (10)
where $(a)$ is by Assumption 2 and Cauchy-Schwartz $(\sum_{j=1}^{N}1\cdot a_{i})^{2}\leq(\sum_{j=1}^{N}a_{i}^{2})\cdot(\sum_{j=1}^{N}1^{2})=N(\sum_{j=1}^{N}a_{i}^{2})$, for real $a_{i}$. $(b)$ is due the independence of noise, $(c)$ is by Assumption 3, and $(d)$ is again by Cauchy-Schwartz. In $(e)$, we let $\mu_{S}=\max_{i}\mathbb{E}[\|f_{i}(0,S_{i,k+1})\|^{2}|\mathcal{H}_{k}]$ and $L_{S}=\max_{i}\mathbb{E}[L_{S_{i,k+1}}^{2}|\mathcal{H}_{k}]$. $(f)$ is due to the normally-distributed channel random variables, $\mathbb{E}[h_{i,k+1}^{2}\sum_{j\neq i}h_{j,k}^{2}]=(N-1)\sigma_{h}^{4}$, $\mathbb{E}[h_{i,k+1}^{2}h_{i,k}^{2}|\mathcal{H}_{k}]=\sigma_{h}^{4}+2K_{hh}^{2}$, and $\mathbb{E}[h_{i,k+1}^{2}\sum_{j<l}h_{j,k}h_{l,k}]$ is equal to zero as one of $h_{j,k}$ and $h_{l,k}$ will always be independent of the other terms in the product and has a mean equal to zero.
@@ B-C Proof of Lemma 4: Norm of the Bias
$\begin{split}&\|b_{k}\|\\ \overset{(a)}{\leq}&\gamma_{k}\frac{\sigma_{h}^{2}}{\alpha_{2}K_{hh}}\sum_{i=1}^{N}\mathbb{E}\Big{[}2N\Big{|}h_{i,k+1}\sum_{j=1}^{N}\Big{(}\frac{h_{j,k}^{2}}{\sigma_{h}^{4}}+n_{j,k}^{2}\Big{)}\Big{|}\times\\ &\hskip 81.09052pt\|\Phi_{k}\|\|\Phi_{k}^{T}\|\|\nabla^{2}\tilde{F}_{i}(\breve{\theta}_{k})\|\|\Phi_{k}\|\Big{|}\mathcal{H}_{k}\Big{]}\\ \leq&2\gamma_{k}\frac{N\alpha_{1}\alpha_{3}^{3}\sigma_{h}^{2}}{\alpha_{2}K_{hh}}\sum_{i=1}^{N}\mathbb{E}\Big{[}\Big{|}h_{i,k+1}\sum_{j=1}^{N}\Big{(}\frac{h_{j,k}^{2}}{\sigma_{h}^{4}}+n_{j,k}^{2}\Big{)}\Big{|}\Big{|}\mathcal{H}_{k}\Big{]}\\ \overset{(b)}{\leq}&2\gamma_{k}\frac{N\alpha_{1}\alpha_{3}^{3}\sigma_{h}^{2}}{\alpha_{2}K_{hh}}\sum_{i=1}^{N}\Big{[}\sigma_{h}\sqrt{\frac{2}{\pi}}\Big{(}2K_{hh}+\sqrt{\sigma_{h}^{4}-K_{hh}^{2}}\Big{)}\\ &\hskip 79.6678pt+(N-1)\sigma_{h}^{3}\sqrt{\frac{2}{\pi}}+N\sqrt{\frac{2}{\pi}}\sigma_{h}\sigma_{n}^{2}\Big{]}\\ =&2\gamma_{k}\frac{N^{2}\alpha_{1}\alpha_{3}^{3}\sigma_{h}^{3}}{\alpha_{2}K_{hh}}\sqrt{\frac{2}{\pi}}\times\\ &\hskip 22.76228pt\Big{[}\Big{(}2K_{hh}+\sqrt{\sigma_{h}^{4}-K_{hh}^{2}}\Big{)}+(N-1)\sigma_{h}^{2}+N\sigma_{n}^{2}\Big{]}\\ \overset{(c)}{=}&c_{3}\gamma_{k},\end{split}$ (11)
where $(a)$ is due to Jensen’s inequality and Cauchy-Schwartz, $(b)$ is by using the half-normal distribution for normal random variables in absolute value explained in the following paragraph, and in $(c)$, $c_{3}=2\frac{N^{2}\alpha_{1}\alpha_{3}^{3}\sigma_{h}^{3}}{\alpha_{2}K_{hh}}\sqrt{\frac{2}{\pi}}\Big{[}\Big{(}2K_{hh}+\sqrt{\sigma_{h}^{4}-K_{hh}^{2}}\Big{)}+(N-1)\sigma_{h}^{2}+N\sigma_{n}^{2}\Big{]}$.
Let $X$ and $Y$ be two variables representing time-correlated channel realizations at times $k$ and $k^{\prime}$ respectively. Assume they follow the $\mathcal{N}(0,\sigma)$ distribution and they have a correlation coefficient $\varrho$. Then, we can write $Y=\varrho X+\sqrt{1-\varrho^{2}}Z$, where $Z$ is independent of X and following the same distribution $\mathcal{N}(0,\sigma)$. Then, $\mathbb{E}[|YX^{2}|]\leq\mathbb{E}[\varrho|X^{3}|+\sqrt{1-\varrho^{2}}|ZX^{2}|]=2\varrho\sqrt{\frac{2}{\pi}}\sigma^{3}+\sqrt{1-\varrho^{2}}\sqrt{\frac{2}{\pi}}\sigma\times\sigma^{2}=(2\varrho+\sqrt{1-\varrho^{2}})\sqrt{\frac{2}{\pi}}\sigma^{3}.$ If we substitute $\sigma=\sigma_{h}$ and $\varrho=\frac{K_{hh}}{\sigma_{h}^{2}}$, we obtain the previous inequality $(b)$.
@@ C-A Stochastic Noise
(ii) $\mathbb{E}[\|X_{K\textquoteright}\|^{2}]<\infty$
@@
In addition, $e_{k}$ and $e_{k^{\prime}}$ are uncorrelated for any $k\neq k^{\prime}$ since (assuming $k>k^{\prime}$) $\mathbb{E}\big{[}e_{k}^{T}e_{k\textquoteright}\big{]}=\mathbb{E}\big{[}\mathbb{E}[e_{k}^{T}e_{k\textquoteright}|\mathcal{H}_{k}]\big{]}=\mathbb{E}\big{[}e_{k\textquoteright}\mathbb{E}[e_{k}^{T}|\mathcal{H}_{k}]\big{]}=0$. Thus,
$\begin{split}\mathbb{E}(\|\sum_{k=K}^{K^{\prime}}\alpha_{k}e_{k}\|^{2})&=\mathbb{E}(\sum_{k=K}^{K^{\prime}}\sum_{k^{\prime}=K}^{K^{\prime}}\alpha_{k}\alpha_{k^{\prime}}\langle e_{k},e_{k^{\prime}}\rangle)\\ &\overset{(a)}{=}\mathbb{E}(\sum_{k=K}^{K^{\prime}}\|\alpha_{k}e_{k}\|^{2})\\ &\leq\sum_{k=K}^{\infty}\mathbb{E}(\alpha_{k}^{2}\|g_{k}-\mathbb{E}[g_{k}|\mathcal{H}_{k}]\|^{2})\\ &=\sum_{k=K}^{\infty}\alpha_{k}^{2}\mathbb{E}(\|g_{k}\|^{2})-\mathbb{E}_{\mathcal{H}_{k}}(\|\mathbb{E}[g_{k}|\mathcal{H}_{k}]\|^{2})\\ &\leq\sum_{k=K}^{\infty}\alpha_{k}^{2}\mathbb{E}(\|g_{k}\|^{2})\overset{(b)}{\leq}c_{2}\sum_{k=K}^{\infty}\alpha_{k}^{2}\gamma_{k}^{2}\overset{(c)}{<}\infty,\\ \end{split}$ (12)
@@
$\begin{split}\mathbb{P}(\sup_{K^{\prime}\geq K}\|\sum_{k=K}^{K^{\prime}}\alpha_{k}e_{k}\|\geq\nu)&\leq\frac{1}{\nu^{2}}\mathbb{E}(\|\sum_{k=K}^{K^{\prime}}\alpha_{k}e_{k}\|^{2})\\ &\overset{(a)}{\leq}\frac{c_{2}}{\nu^{2}}\sum_{k=K}^{\infty}\alpha_{k}^{2}\gamma_{k}^{2},\\ \end{split}$ (13)
@@
Since $c_{2}$ is a bounded constant and $\lim_{K\rightarrow\infty}\sum_{k=K}^{\infty}\alpha_{k}^{2}\gamma_{k}^{2}=0$ by Assumption 5, we get $\lim_{K\rightarrow\infty}\frac{c_{2}}{\nu^{2}}\sum_{k=K}^{\infty}\alpha_{k}^{2}\gamma_{k}^{2}=0$ for any bounded constant $\nu$. Hence, the probability that $\|\sum_{k=K}^{K^{\prime}}\alpha_{k}e_{k}\|\geq\nu$ also vanishes as $K\rightarrow\infty$, which concludes the proof.
@@ C-B Proof of Theorem 1: Convergence Analysis
By the $L$-smoothness assumption and the algorithm update step $\theta_{k+1}=\theta_{k}-\alpha_{k}g_{k}$, we have
$\begin{split}&F(\theta_{k+1})\\ \leq&F(\theta_{k})-\alpha_{k}\langle\nabla F(\theta_{k}),g_{k}\rangle+\frac{\alpha_{k}^{2}L}{2}\|g_{k}\|^{2}\\ =&F(\theta_{k})-\alpha_{k}\langle\nabla F(\theta_{k}),e_{k}+\mathbb{E}[g_{k}|\mathcal{H}_{k}]\rangle+\frac{\alpha_{k}^{2}L}{2}\|g_{k}\|^{2}\\ =&F(\theta_{k})-\alpha_{k}\langle\nabla F(\theta_{k}),e_{k}\rangle-c_{1}\alpha_{k}\gamma_{k}\|\nabla F(\theta_{k})\|^{2}\\ &-c_{1}\alpha_{k}\gamma_{k}\langle\nabla F(\theta_{k}),b_{k}\rangle+\frac{\alpha_{k}^{2}L}{2}\|g_{k}\|^{2}\\ \overset{(a)}{\leq}&F(\theta_{k})-\alpha_{k}\langle\nabla F(\theta_{k}),e_{k}\rangle-c_{1}\alpha_{k}\gamma_{k}\|\nabla F(\theta_{k})\|^{2}\\ &+\frac{c_{1}\alpha_{k}\gamma_{k}}{2}\|\nabla F(\theta_{k})\|^{2}+\frac{c_{1}\alpha_{k}\gamma_{k}}{2}\|b_{k}\|^{2}+\frac{\alpha_{k}^{2}L}{2}\|g_{k}\|^{2}\\ =&F(\theta_{k})-\alpha_{k}\langle\nabla F(\theta_{k}),e_{k}\rangle-\frac{c_{1}\alpha_{k}\gamma_{k}}{2}\|\nabla F(\theta_{k})\|^{2}\\ &+\frac{c_{1}\alpha_{k}\gamma_{k}}{2}\|b_{k}\|^{2}+\frac{\alpha_{k}^{2}L}{2}\|g_{k}\|^{2}\\ \end{split}$ (14)
where $(a)$ is by $-\langle a,b\rangle\leq\frac{1}{2}\|a\|^{2}+\frac{1}{2}\|b\|^{2}$.
@@
$\begin{split}F(\theta^{*})\leq F(\theta_{K+1})\leq&F(\theta_{0})-\frac{c_{1}}{2}\sum_{k=0}^{K}\alpha_{k}\gamma_{k}\|\nabla F(\theta_{k})\|^{2}\\ &-\sum_{k=0}^{K}\alpha_{k}\langle\nabla F(\theta_{k}),e_{k}\rangle\\ &+\frac{c_{1}}{2}\sum_{k=0}^{K}\alpha_{k}\gamma_{k}\|b_{k}\|^{2}+\frac{c_{2}L}{2}\sum_{k=0}^{K}\alpha_{k}^{2}\|g_{k}\|^{2}\\ \end{split}$ (15)
Hence,
$\begin{split}\sum_{k=0}^{K}\alpha_{k}\gamma_{k}\|\nabla F(\theta_{k})\|^{2}\leq&\frac{2}{c_{1}}(F(\theta_{0})-F(\theta^{*}))\\ &-\frac{2}{c_{1}}\sum_{k=0}^{K}\alpha_{k}\langle\nabla F(\theta_{k}),e_{k}\rangle\\ &+\sum_{k=0}^{K}\alpha_{k}\gamma_{k}\|b_{k}\|^{2}+\frac{c_{2}L}{c_{1}}\sum_{k=0}^{K}\alpha_{k}^{2}\|g_{k}\|^{2}\\ \end{split}$ (16)
By Assumption 3, $\|\nabla F(\theta_{k})\|$ is bounded for any $\theta_{k}\in\mathbb{R}^{d}$ and by taking the summation in (13) between 00 and $\infty$, we have
$\lim_{K\rightarrow\infty}\|\sum_{k=0}^{K}\alpha_{k}\langle\nabla F(\theta_{k}),e_{k}\rangle\|<\infty.$ (17)
From Lemma 4, we know that $\|b_{k}\|^{2}\sim\gamma_{k}^{2}$. Hence, by Assumption 5, $\lim_{K\rightarrow\infty}\sum_{k=0}^{K}\alpha_{k}\gamma_{k}^{3}<\infty$.
From Lemma 2 and by looking closely at the use of the Lipschitz continuity property in (10), we can say $\|g_{k}\|^{2}\leq c$ for some $c>0$. Thus, again by Assumption 5, $\lim_{K\rightarrow\infty}\sum_{k=0}^{K}\alpha_{k}^{2}<\infty.$ We conclude that
@@
$\begin{split}\|\nabla F(\theta_{k_{l}+1})\|&\geq\|\nabla F(\theta_{k_{l}})\|-\|\nabla F(\theta_{k_{l}+1})-\nabla F(\theta_{k_{l}})\|\\ &\geq\rho-\varepsilon-L\|\theta_{k_{l}+1}-\theta_{k_{l}}\|\\ &=\rho-\varepsilon-L\alpha_{k_{l}}\|g_{k_{l}}\|\\ &\geq\rho-\varepsilon-L\sqrt{c}\alpha_{k_{l}},\end{split}$ (19)

### v2

@@ B-B Proof of Lemma 2: Expected Norm Squared of the Estimated Gradient
$\begin{split}&\mathbb{E}[\|g_{k}\|^{2}|\mathcal{H}_{k}]\\ =&\mathbb{E}\Big{[}\|\Phi_{k}\|^{2}\Big{(}\sum_{i=1}^{N}h_{i,k+1}\tilde{f}_{i}\big{(}\theta_{k}^{\prime},S_{i,k+1}\big{)}+n_{i,k+1}\Big{)}^{2}\Big{|}\mathcal{H}_{k}\Big{]}\\ \overset{(a)}{\leq}&\beta_{2}^{2}\mathbb{E}\Big{[}\Big{(}\sum_{i=1}^{N}h_{i,k+1}\tilde{f}_{i}\big{(}\theta_{k}^{\prime},S_{i,k+1}\big{)}+n_{i,k+1}\Big{)}^{2}\Big{|}\mathcal{H}_{k}\Big{]}\\ =&\beta_{2}^{2}\mathbb{E}\Big{[}\Big{(}\sum_{i=1}^{N}h_{i,k+1}\tilde{f}_{i}\big{(}\theta_{k}^{\prime},S_{i,k+1}\big{)}\Big{)}^{2}+\Big{(}\sum_{i=1}^{N}n_{i,k+1}\Big{)}^{2}\\ &\hskip 28.45274pt+2\Big{(}\sum_{i=1}^{N}h_{i,k+1}\tilde{f}_{i}\big{(}\theta_{k}^{\prime},S_{i,k+1}\big{)}\Big{)}\Big{(}\sum_{i=1}^{N}n_{i,k+1}\Big{)}\Big{|}\mathcal{H}_{k}\Big{]}\\ \overset{(b)}{=}&\beta_{2}^{2}\mathbb{E}\Big{[}\Big{(}\sum_{i=1}^{N}h_{i,k+1}\tilde{f}_{i}\big{(}\theta_{k}^{\prime},S_{i,k+1}\big{)}\Big{)}^{2}+\Big{(}\sum_{i=1}^{N}n_{i,k+1}\Big{)}^{2}\Big{|}\mathcal{H}_{k}\Big{]}\\ \overset{(c)}{=}&\beta_{2}^{2}\mathbb{E}\Big{[}N\sum_{i=1}^{N}\Big{(}h_{i,k+1}^{2}\frac{C^{2}}{\sigma_{h}^{4}}\Big{)}+\sum_{i=1}^{N}n_{i,k+1}^{2}\\ &\hskip 28.45274pt+2\sum_{j<i}n_{i,k+1}n_{j,k+1}\Big{|}\mathcal{H}_{k}\Big{]}\\ \overset{(d)}{=}&\beta_{2}^{2}\sum_{i=1}^{N}\mathbb{E}\Big{[}Nh_{i,k+1}^{2}\frac{C^{2}}{\sigma_{h}^{4}}+n_{i,k+1}^{2}\Big{|}\mathcal{H}_{k}\Big{]}\\ =&\beta_{2}^{2}N\Big{(}\frac{NC^{2}}{\sigma_{h}^{2}}+\sigma_{n}^{2}\Big{)}\\ :=&c_{2},\end{split}$ (9)
where $(a)$ is by Assumption 2, $(b)$ is due to the due to the independence of $n_{i,k+1}$ and its zero-mean, $(c)$ is by Cauchy-Schwartz $(\sum_{j=1}^{N}1\cdot a_{i})^{2}\leq(\sum_{j=1}^{N}a_{i}^{2})\cdot(\sum_{j=1}^{N}1^{2})=N(\sum_{j=1}^{N}a_{i}^{2})$, for real $a_{i}$, and Assumption 3. $(d)$ is due to the uncorrelatedness of the noise.
@@ B-C Proof of Lemma 4: Norm of the Bias
$\begin{split}\|b_{k}\|\overset{(a)}{\leq}&\gamma_{k}\frac{\sigma_{h}^{2}}{\beta_{1}K_{hh}}\sum_{i=1}^{N}\mathbb{E}\Big{[}2N\Big{|}h_{i,k+1}\sum_{j=1}^{N}\Big{(}\frac{h_{j,k}^{2}}{\sigma_{h}^{4}}+n_{j,k}^{2}\Big{)}\Big{|}\times\\ &\hskip 81.09052pt\|\Phi_{k}\|\|\Phi_{k}^{T}\|\|\nabla^{2}\tilde{F}_{i}(\breve{\theta}_{k})\|\|\Phi_{k}\|\Big{|}\mathcal{H}_{k}\Big{]}\\ \leq&2\gamma_{k}\frac{N\beta_{3}\beta_{2}^{3}\sigma_{h}^{2}}{\beta_{1}K_{hh}}\sum_{i=1}^{N}\mathbb{E}\Big{[}\Big{|}h_{i,k+1}\sum_{j=1}^{N}\Big{(}\frac{h_{j,k}^{2}}{\sigma_{h}^{4}}+n_{j,k}^{2}\Big{)}\Big{|}\Big{|}\mathcal{H}_{k}\Big{]}\\ \end{split}$
$\begin{split}\overset{(b)}{\leq}&2\gamma_{k}\frac{N\beta_{3}\beta_{2}^{3}\sigma_{h}^{2}}{\beta_{1}K_{hh}}\sum_{i=1}^{N}\Big{[}\sigma_{h}\sqrt{\frac{2}{\pi}}\Big{(}2K_{hh}+\sqrt{\sigma_{h}^{4}-K_{hh}^{2}}\Big{)}\\ &\hskip 79.6678pt+(N-1)\sigma_{h}^{3}\sqrt{\frac{2}{\pi}}+N\sqrt{\frac{2}{\pi}}\sigma_{h}\sigma_{n}^{2}\Big{]}\\ =&2\gamma_{k}\frac{N^{2}\beta_{3}\beta_{2}^{3}\sigma_{h}^{3}}{\beta_{1}K_{hh}}\sqrt{\frac{2}{\pi}}\times\\ &\hskip 22.76228pt\Big{[}\Big{(}2K_{hh}+\sqrt{\sigma_{h}^{4}-K_{hh}^{2}}\Big{)}+(N-1)\sigma_{h}^{2}+N\sigma_{n}^{2}\Big{]}\\ :=&c_{3}\gamma_{k},\end{split}$ (10)
where $(a)$ is due to Jensen’s inequality and Cauchy-Schwartz. $(b)$ is by using the half-normal distribution for normal random variables in absolute value: Let $X$ and $Y$ be two variables representing time-correlated channel realizations at times $k$ and $k^{\prime}$ respectively. Assume they follow the $\mathcal{N}(0,\sigma)$ distribution and they have a correlation coefficient $\varrho$. Then, we can write $Y=\varrho X+\sqrt{1-\varrho^{2}}Z$, where $Z$ is independent of X and following the same distribution $\mathcal{N}(0,\sigma)$. Then, $\mathbb{E}[|YX^{2}|]\leq\mathbb{E}[\varrho|X^{3}|+\sqrt{1-\varrho^{2}}|ZX^{2}|]=2\varrho\sqrt{\frac{2}{\pi}}\sigma^{3}+\sqrt{1-\varrho^{2}}\sqrt{\frac{2}{\pi}}\sigma\times\sigma^{2}=(2\varrho+\sqrt{1-\varrho^{2}})\sqrt{\frac{2}{\pi}}\sigma^{3}.$ If we substitute $\sigma=\sigma_{h}$ and $\varrho=\frac{K_{hh}}{\sigma_{h}^{2}}$, we obtain the previous inequality $(b)$.
@@ C-A Stochastic Noise
(ii) $\mathbb{E}[\|X_{K\textquoteright}\|^{2}]<\infty$.
@@
In addition, $e_{k}$ and $e_{k^{\prime}}$ are uncorrelated for any $k\neq k^{\prime}$ since (assuming $k>k^{\prime}$) $\mathbb{E}\big{[}e_{k}^{T}e_{k\textquoteright}\big{]}=\mathbb{E}\big{[}\mathbb{E}[e_{k}^{T}e_{k\textquoteright}|\mathcal{H}_{k}]\big{]}=\mathbb{E}\big{[}e_{k\textquoteright}\mathbb{E}[e_{k}^{T}|\mathcal{H}_{k}]\big{]}=0$.
Thus,
$\begin{split}\mathbb{E}(\|\sum_{k=K}^{K^{\prime}}\alpha_{k}e_{k}\|^{2})&=\mathbb{E}(\sum_{k=K}^{K^{\prime}}\sum_{k^{\prime}=K}^{K^{\prime}}\alpha_{k}\alpha_{k^{\prime}}\langle e_{k},e_{k^{\prime}}\rangle)\\ &\overset{(a)}{=}\mathbb{E}(\sum_{k=K}^{K^{\prime}}\|\alpha_{k}e_{k}\|^{2})\\ &\leq\sum_{k=K}^{\infty}\mathbb{E}(\alpha_{k}^{2}\|g_{k}-\mathbb{E}[g_{k}|\mathcal{H}_{k}]\|^{2})\\ &=\sum_{k=K}^{\infty}\alpha_{k}^{2}\mathbb{E}(\|g_{k}\|^{2})-\mathbb{E}_{\mathcal{H}_{k}}(\|\mathbb{E}[g_{k}|\mathcal{H}_{k}]\|^{2})\\ &\leq\sum_{k=K}^{\infty}\alpha_{k}^{2}\mathbb{E}(\|g_{k}\|^{2})\overset{(b)}{\leq}c_{2}\sum_{k=K}^{\infty}\alpha_{k}^{2}\overset{(c)}{<}\infty,\\ \end{split}$ (11)
@@
$\begin{split}\mathbb{P}(\sup_{K^{\prime}\geq K}\|\sum_{k=K}^{K^{\prime}}\alpha_{k}e_{k}\|\geq\nu)&\leq\frac{1}{\nu^{2}}\mathbb{E}(\|\sum_{k=K}^{K^{\prime}}\alpha_{k}e_{k}\|^{2})\\ &\overset{(a)}{\leq}\frac{c_{2}}{\nu^{2}}\sum_{k=K}^{\infty}\alpha_{k}^{2},\end{split}$ (12)
@@
Since $c_{2}$ is a bounded constant and $\lim_{K\rightarrow\infty}\sum_{k=K}^{\infty}\alpha_{k}^{2}=0$ by Assumption 5, we get $\lim_{K\rightarrow\infty}\frac{c_{2}}{\nu^{2}}\sum_{k=K}^{\infty}\alpha_{k}^{2}=0$ for any bounded constant $\nu$. Hence, the probability that $\|\sum_{k=K}^{K^{\prime}}\alpha_{k}e_{k}\|\geq\nu$ also vanishes as $K\rightarrow\infty$, which concludes the proof.
@@ C-B Proof of Theorem 1: Convergence Analysis
By the $L$-smoothness assumption, the algorithm update step $\theta_{k+1}=\theta_{k}-\alpha_{k}g_{k}$, and knowing $g_{k}=e_{k}+\mathbb{E}[g_{k}|\mathcal{H}_{k}]$, we have
$\begin{split}F(\theta_{k+1})\leq&F(\theta_{k})-\alpha_{k}\langle\nabla F(\theta_{k}),g_{k}\rangle+\frac{\alpha_{k}^{2}L}{2}\|g_{k}\|^{2}\\ =&F(\theta_{k})-\alpha_{k}\langle\nabla F(\theta_{k}),e_{k}\rangle-c_{1}\alpha_{k}\gamma_{k}\|\nabla F(\theta_{k})\|^{2}\\ &-c_{1}\alpha_{k}\gamma_{k}\langle\nabla F(\theta_{k}),b_{k}\rangle+\frac{\alpha_{k}^{2}L}{2}\|g_{k}\|^{2}\\ \overset{(a)}{\leq}&F(\theta_{k})-\alpha_{k}\langle\nabla F(\theta_{k}),e_{k}\rangle-c_{1}\alpha_{k}\gamma_{k}\|\nabla F(\theta_{k})\|^{2}\\ +&\frac{c_{1}\alpha_{k}\gamma_{k}}{2}\|\nabla F(\theta_{k})\|^{2}+\frac{c_{1}\alpha_{k}\gamma_{k}}{2}\|b_{k}\|^{2}+\frac{\alpha_{k}^{2}L}{2}\|g_{k}\|^{2}\\ \end{split}$
$\begin{split}\overset{(b)}{\leq}&F(\theta_{k})-\alpha_{k}\langle\nabla F(\theta_{k}),e_{k}\rangle-\frac{c_{1}\alpha_{k}\gamma_{k}}{2}\|\nabla F(\theta_{k})\|^{2}\\ &+\frac{c_{1}c_{3}^{2}\alpha_{k}\gamma_{k}^{3}}{2}+\frac{\alpha_{k}^{2}L}{2}\|g_{k}\|^{2}\end{split}$ (13)
where $(a)$ is by $-\langle a,b\rangle\leq\frac{1}{2}\|a\|^{2}+\frac{1}{2}\|b\|^{2}$ and $(b)$ is by Lemma 4.
@@
$\begin{split}&F(\theta^{*})\leq F(\theta_{K+1})\leq F(\theta_{0})-\frac{c_{1}}{2}\sum_{k=0}^{K}\alpha_{k}\gamma_{k}\|\nabla F(\theta_{k})\|^{2}\\ &-\sum_{k=0}^{K}\alpha_{k}\langle\nabla F(\theta_{k}),e_{k}\rangle+\frac{c_{1}c_{3}^{2}}{2}\sum_{k=0}^{K}\alpha_{k}\gamma_{k}^{3}+\frac{c_{2}L}{2}\sum_{k=0}^{K}\alpha_{k}^{2}\|g_{k}\|^{2}\\ \end{split}$
Hence, knowing that $\delta_{0}=F(\theta_{0})-F(\theta^{*})$,
$\begin{split}\sum_{k=0}^{K}\alpha_{k}\gamma_{k}\|\nabla F(\theta_{k})\|^{2}\leq&\frac{2\delta_{0}}{c_{1}}-\frac{2}{c_{1}}\sum_{k=0}^{K}\alpha_{k}\langle\nabla F(\theta_{k}),e_{k}\rangle\\ &+c_{3}^{2}\sum_{k=0}^{K}\alpha_{k}\gamma_{k}^{3}+\frac{c_{2}L}{c_{1}}\sum_{k=0}^{K}\alpha_{k}^{2}\|g_{k}\|^{2}\\ \end{split}$ (14)
By Assumption 3, $\|\nabla F(\theta_{k})\|$ is bounded for any $\theta_{k}\in\mathbb{R}^{d}$ and by taking the summation in (12) between 00 and $\infty$, we have $\lim_{K\rightarrow\infty}\|\sum_{k=0}^{K}\alpha_{k}\langle\nabla F(\theta_{k}),e_{k}\rangle\|<\infty$. By Assumption 5, $\lim_{K\rightarrow\infty}\sum_{k=0}^{K}\alpha_{k}\gamma_{k}^{3}<\infty$.
From Lemma 2 and by looking closely at the use of the local functions’ boundedness property in (9), we can say $\|g_{k}\|^{2}\leq c$ almost surely for some $c>0$. Thus, again by Assumption 5, $\lim_{K\rightarrow\infty}\sum_{k=0}^{K}\alpha_{k}^{2}<\infty.$ We conclude that
@@
$\begin{split}\|\nabla F(\theta_{k_{l}+1})\|&\geq\|\nabla F(\theta_{k_{l}})\|-\|\nabla F(\theta_{k_{l}+1})-\nabla F(\theta_{k_{l}})\|\\ &\geq\rho-\varepsilon-L\|\theta_{k_{l}+1}-\theta_{k_{l}}\|\\ &=\rho-\varepsilon-L\alpha_{k_{l}}\|g_{k_{l}}\|\\ &\geq\rho-\varepsilon-L\sqrt{c}\alpha_{k_{l}},\end{split}$